# Optimization and Generalization Guarantees for Weight Normalization

### v1

@@ C.1 Restricted Strong Convexity and Smoothness
$(\theta^{\prime}-\bar{\theta})^{\top}\frac{\partial^{2}{\cal L}(\tilde{\theta})}{\partial\theta^{2}}(\theta^{\prime}-\bar{\theta})$
$=\underbrace{\frac{1}{n}\sum_{i=1}^{n}\ell^{\prime\prime}_{i}\left\langle\theta^{\prime}-\bar{\theta},\frac{\partial f(\tilde{\theta};\mathbf{x}_{i})}{\partial\theta}\right\rangle^{2}}_{I_{1}}+\underbrace{\frac{1}{n}\sum_{i=1}^{n}\ell^{\prime}_{i}(\theta^{\prime}-\bar{\theta})^{\top}\frac{\partial^{2}f(\tilde{\theta};\mathbf{x}_{i})}{\partial\theta^{2}}(\theta^{\prime}-\bar{\theta})}_{I_{2}}~{},$
where $\ell_{i}=\ell(y_{i},f(\tilde{\theta},\mathbf{x}_{i})$, $\ell^{\prime}_{i}=\frac{\partial\ell(y_{i},z)}{\partial z}\bigg{\rvert}_{z=f(\tilde{\theta},\mathbf{x}_{i})}$, and $\ell^{\prime\prime}_{i}=\frac{\partial^{2}\ell(y_{i},z)}{\partial z^{2}}\bigg{\rvert}_{z=f(\tilde{\theta},\mathbf{x}_{i})}$. It is easy to note that $I_{1}\leq 2\varrho^{2}_{\theta}\|\theta^{\prime}-\bar{\theta}\|_{2}^{2}$. For $I_{2}$, we have that
$I_{2}$ $=\frac{1}{n}\sum_{i=1}^{n}\ell^{\prime}_{i}(\theta^{\prime}-\bar{\theta})^{\top}\frac{\partial^{2}f(\tilde{\theta};\mathbf{x}_{i})}{\partial\theta^{2}}(\theta^{\prime}-\bar{\theta})$
$\leq\left|\sum_{i=1}^{n}\left(\frac{1}{\sqrt{n}}\ell^{\prime}_{i}\right)\left(\frac{1}{\sqrt{n}}\|\theta^{\prime}-\bar{\theta}\|_{2}^{2}\left\|\frac{\partial^{2}f(\tilde{\theta};\mathbf{x}_{i})}{\partial\theta^{2}}\right\|_{2}\right)\right|$
$\overset{(a)}{\leq}\left(\frac{1}{n}\|[\ell^{\prime}_{i}]_{i}\|_{2}^{2}\right)^{1/2}\left(\frac{1}{n}\sum_{i=1}^{n}\left\lVert\theta^{\prime}-\bar{\theta}\right\rVert_{2}^{4}\left\lVert\frac{\partial^{2}f(\tilde{\theta};\mathbf{x}_{i})}{\partial\theta^{2}}\right\rVert_{2}^{2}\right)^{1/2}$
$\overset{(b)}{\leq}2\sqrt{{\cal L}(\tilde{\theta})}\left(\frac{1}{n}\sum_{i=1}^{n}\left\lVert\frac{\partial^{2}f(\tilde{\theta};\mathbf{x}_{i})}{\partial\theta^{2}}\right\rVert_{2}^{2}\right)^{1/2}\left\lVert\theta^{\prime}-\bar{\theta}\right\rVert_{2}^{2}~{},$
@@
$(\theta^{\prime}-\bar{\theta})^{\top}\frac{\partial^{2}{\cal L}(\tilde{\theta})}{\partial\theta^{2}}(\theta^{\prime}-\bar{\theta})$ $\leq\left[2\varrho^{2}_{\theta}+2\sqrt{\varphi}\left(\frac{1}{n}\sum_{i=1}^{n}\left\lVert\frac{\partial^{2}f(\tilde{\theta};\mathbf{x}_{i})}{\partial\theta^{2}}\right\rVert_{2}^{2}\right)^{1/2}\right]\left\lVert\theta^{\prime}-\bar{\theta}\right\rVert_{2}^{2}~{}.$
@@
$\varrho^{2}_{\theta}\leq{\cal O}\left((1+L^{2}|\phi(0)|^{2}m)\left(1+\frac{L(1+\rho_{1})^{2}}{{\min_{\begin{subarray}{c}i\in[m]\\ l\in[L]\end{subarray}}\left\lVert W^{(l)}_{i}\right\rVert_{2}^{2}}}\right)\right)$
@@
$2\varrho^{2}_{\theta}+2\sqrt{\varphi}\left(\frac{1}{n}\sum_{i=1}^{n}\left\lVert\frac{\partial^{2}f(\tilde{\theta};\mathbf{x}_{i})}{\partial\theta^{2}}\right\rVert_{2}^{2}\right)^{1/2}$
$\qquad+{\cal O}\left(\frac{(1+\rho_{1})(1+L|\phi(0)|\sqrt{m})\cdot(1+\rho_{1})L^{3}(\frac{1}{\sqrt{m}}+L^{2}\max\{|\phi(0)|,|\phi(0)|^{2}\})}{\min\{{\min_{\begin{subarray}{c}i\in[m]\\ l\in[L]\end{subarray}}\left\lVert W^{(l)}_{i}\right\rVert_{2}},{\min_{\begin{subarray}{c}i\in[m]\\ l\in[L]\end{subarray}}\left\lVert W^{(l)}_{i}\right\rVert_{2}^{2}}\}}\right)$
$\leq{\cal O}\left(L^{3}(1+L^{2}\max\{|\phi(0)|,|\phi(0)|^{2}\}m)(1+\rho_{1})^{2}\left(1+\frac{L}{{\min_{\begin{subarray}{c}i\in[m]\\ l\in[L]\end{subarray}}\left\lVert W^{(l)}_{i}\right\rVert_{2}^{2}}}\right.\right.$
$\left.\left.\qquad+\frac{(\frac{1}{\sqrt{m}}+L^{2}\max\{|\phi(0)|,|\phi(0)|^{2}\})}{\min\{{\min_{\begin{subarray}{c}i\in[m]\\ l\in[L]\end{subarray}}\left\lVert W^{(l)}_{i}\right\rVert_{2}},{\min_{\begin{subarray}{c}i\in[m]\\ l\in[L]\end{subarray}}\left\lVert W^{(l)}_{i}\right\rVert_{2}^{2}}\}}\right)\right)$
$\leq{\cal O}\left(L^{4}(1+L^{2}\max\{|\phi(0)|,|\phi(0)|^{2}\}m)^{2}(1+\rho_{1})^{2}\left(1+\frac{1}{\min\{{\min_{\begin{subarray}{c}i\in[m]\\ l\in[L]\end{subarray}}\left\lVert W^{(l)}_{i}\right\rVert_{2}},{\min_{\begin{subarray}{c}i\in[m]\\ l\in[L]\end{subarray}}\left\lVert W^{(l)}_{i}\right\rVert_{2}^{2}}\}}\right)\right).$
@@
Appendix D Generalization bound for WeightNorm

### v2

@@ C.1 Restricted Strong Convexity and Smoothness
where $\ell_{i}=\ell(y_{i},f(\tilde{\theta},\mathbf{x}_{i})$, $\ell^{\prime}_{i}=\frac{\partial\ell(y_{i},z)}{\partial z}\bigg{\rvert}_{z=f(\tilde{\theta},\mathbf{x}_{i})}$, and $\ell^{\prime\prime}_{i}=\frac{\partial^{2}\ell(y_{i},z)}{\partial z^{2}}\bigg{\rvert}_{z=f(\tilde{\theta},\mathbf{x}_{i})}$. It is easy to note that $I_{1}\leq 2\varrho^{2}_{\tilde{\theta}}\|\theta^{\prime}-\theta\|_{2}^{2}$. For $I_{2}$, we have that
$\leq\left|\sum_{i=1}^{n}\left(\frac{1}{\sqrt{n}}\ell^{\prime}_{i}\right)\left(\frac{1}{\sqrt{n}}\|\theta^{\prime}-\theta\|_{2}^{2}\left\|\frac{\partial^{2}f(\tilde{\theta};\mathbf{x}_{i})}{\partial\theta^{2}}\right\|_{2}\right)\right|$
$\overset{(a)}{\leq}\left(\frac{1}{n}\|[\ell^{\prime}_{i}]_{i}\|_{2}^{2}\right)^{1/2}\left(\frac{1}{n}\sum_{i=1}^{n}\left\lVert\theta^{\prime}-\theta\right\rVert_{2}^{4}\left\lVert\frac{\partial^{2}f(\tilde{\theta};\mathbf{x}_{i})}{\partial\theta^{2}}\right\rVert_{2}^{2}\right)^{1/2}$
$\overset{(b)}{\leq}2\sqrt{{\cal L}(\tilde{\theta})}\left(\frac{1}{n}\sum_{i=1}^{n}\left\lVert\frac{\partial^{2}f(\tilde{\theta};\mathbf{x}_{i})}{\partial\theta^{2}}\right\rVert_{2}^{2}\right)^{1/2}\left\lVert\theta^{\prime}-\theta\right\rVert_{2}^{2}~{},$
@@
$(\theta^{\prime}-\theta)^{\top}\frac{\partial^{2}{\cal L}(\tilde{\theta})}{\partial\theta^{2}}(\theta^{\prime}-\theta)$ $\leq\left[2\varrho^{2}_{\tilde{\theta}}+2\sqrt{\varphi}\left(\frac{1}{n}\sum_{i=1}^{n}\left\lVert\frac{\partial^{2}f(\tilde{\theta};\mathbf{x}_{i})}{\partial\theta^{2}}\right\rVert_{2}^{2}\right)^{1/2}\right]\left\lVert\theta^{\prime}-\theta\right\rVert_{2}^{2}~{}.$
@@
$\varrho^{2}_{\tilde{\theta}}\leq{\cal O}\left((1+L^{2}|\phi(0)|^{2}m)\left(1+\frac{L(1+\rho_{1})^{2}}{\min_{\begin{subarray}{c}i\in[m]\\ l\in[L]\end{subarray}}\left\lVert\tilde{W}_{i}^{(l)}\right\rVert_{2}^{2}}\right)\right)$
@@
$2\varrho^{2}_{\tilde{\theta}}+2\sqrt{\varphi}\left(\frac{1}{n}\sum_{i=1}^{n}\left\lVert\frac{\partial^{2}f(\tilde{\theta};\mathbf{x}_{i})}{\partial\theta^{2}}\right\rVert_{2}^{2}\right)^{1/2}$
$\leq{\cal O}\left((1+L^{2}|\phi(0)|^{2}m)\left(1+\frac{L(1+\rho_{1})^{2}}{\min_{\begin{subarray}{c}i\in[m]\\ l\in[L]\end{subarray}}\left\lVert\tilde{W}_{i}^{(l)}\right\rVert_{2}^{2}}\right)\right)$
$\qquad+{\cal O}\left(\frac{(1+\rho_{1})(1+L|\phi(0)|\sqrt{m})\cdot(1+\rho_{1})L^{3}(\frac{1}{\sqrt{m}}+L^{2}\max\{|\phi(0)|,|\phi(0)|^{2}\})}{\min\{\min_{\begin{subarray}{c}i\in[m]\\ l\in[L]\end{subarray}}\left\lVert\tilde{W}_{i}^{(l)}\right\rVert_{2},\min_{\begin{subarray}{c}i\in[m]\\ l\in[L]\end{subarray}}\left\lVert\tilde{W}_{i}^{(l)}\right\rVert_{2}^{2}\}}\right)$
$\leq{\cal O}\left(L^{3}(1+L^{2}\max\{|\phi(0)|,|\phi(0)|^{2}\}m)(1+\rho_{1})^{2}\left(1+\frac{L}{\min_{\begin{subarray}{c}i\in[m]\\ l\in[L]\end{subarray}}\left\lVert\tilde{W}_{i}^{(l)}\right\rVert_{2}^{2}}\right.\right.$
$\left.\left.\qquad+\frac{(\frac{1}{\sqrt{m}}+L^{2}\max\{|\phi(0)|,|\phi(0)|^{2}\})}{\min\{\min_{\begin{subarray}{c}i\in[m]\\ l\in[L]\end{subarray}}\left\lVert\tilde{W}_{i}^{(l)}\right\rVert_{2},\min_{\begin{subarray}{c}i\in[m]\\ l\in[L]\end{subarray}}\left\lVert\tilde{W}_{i}^{(l)}\right\rVert_{2}^{2}\}}\right)\right)$
$\leq{\cal O}\left(L^{4}(1+L^{2}\max\{|\phi(0)|,|\phi(0)|^{2}\}m)^{2}(1+\rho_{1})^{2}\left(1+\frac{1}{\min\{\min_{\begin{subarray}{c}i\in[m]\\ l\in[L]\end{subarray}}\left\lVert\tilde{W}_{i}^{(l)}\right\rVert_{2},\min_{\begin{subarray}{c}i\in[m]\\ l\in[L]\end{subarray}}\left\lVert\tilde{W}_{i}^{(l)}\right\rVert_{2}^{2}\}}\right)\right).$
@@
C.2 About the adaptation of [Banerjee et al., 2023, Theorem 5.3]
To adapt [Banerjee et al., 2023, Theorem 5.3] to our setting, we need to adapt the supporting lemma [Banerjee et al., 2023, Lemma 5.1] to our setting as follows:
Lemma C.1 (Auxiliary result using RSC, [Banerjee et al., 2023, Lemma 5.1]).
Let $B_{t}:=Q^{\theta_{t}}_{\kappa}\cap B_{\rho_{2}}^{\operatorname{Euc}}(\theta_{t})\cap\{\theta\in\mathbb{R}^{p}\,|\,\mathbf{v}\in B_{\rho_{1}}^{\operatorname{Euc}}(\mathbf{v}_{0})\}$ and $\overline{\theta}_{t}\in\operatorname{arginf}_{\theta\in B_{t}}{\cal L}(\theta)$. Using the setting of Lemma 5.1, if $\alpha_{\theta_{t},\overline{\theta}_{t}}>0$, then
${\cal L}(\theta_{t})-\inf_{\theta\in B_{t}}{\cal L}(\theta)\leq\frac{1}{2\alpha_{\theta_{t},\overline{\theta}_{t}}}\|\nabla_{\theta}{\cal L}(\theta_{t})\|_{2}^{2}~{}.$ (41)
Proof.
By Lemma 5.1 we have
${\cal L}(\overline{\theta}_{t})\geq{\cal L}(\theta_{t})+\langle\overline{\theta}_{t}-\theta_{t},\nabla_{\theta}{\cal L}(\theta_{t})\rangle+\frac{\alpha_{\theta_{t},\overline{\theta}_{t}}}{2}\|\overline{\theta}_{t}-\theta_{t}\|_{2}^{2}~{}.$ (42)
Now, let us take $\alpha_{\theta_{t},\overline{\theta}_{t}}$ and define
^$\hat{{\cal L}}_{t}(\theta):={\cal L}(\theta_{t})+\langle\theta-\theta_{t},\nabla_{\theta}{\cal L}(\theta_{t})\rangle+\frac{\alpha_{\theta_{t},\overline{\theta}_{t}}}{2}\|\theta-\theta_{t}\|_{2}^{2}~{}.$
Note that ^$\hat{{\cal L}}_{t}(\theta)$ is minimized at $\hat{\theta}_{t+1}:=\theta_{t}-\nabla_{\theta}{\cal L}(\theta_{t})/\alpha_{\theta_{t},\overline{\theta}_{t}}$ and the minimum value is:
$\inf_{\theta\in\mathbb{R}^{p}}\hat{{\cal L}}_{\theta_{t}}(\theta)=\hat{{\cal L}}_{t}(\hat{\theta}_{t+1})={\cal L}(\theta_{t})-\frac{1}{2\alpha_{\theta_{t},\overline{\theta}_{t}}}\|\nabla_{\theta}{\cal L}(\theta_{t})\|_{2}^{2}~{}.$
Then, we have
$\inf_{\theta\in B_{t}}{\cal L}(\theta)={\cal L}(\overline{\theta}_{t})\overset{(a)}{\geq}\hat{{\cal L}}_{t}(\overline{\theta}_{t})\geq\inf_{\theta\in\mathbb{R}^{p}}\hat{{\cal L}}_{t}(\theta)={\cal L}(\theta_{t})-\frac{1}{2\alpha_{\theta_{t},\overline{\theta}_{t}}}\|\nabla_{\theta}{\cal L}(\theta_{t})\|_{2}^{2}~{},$
where (a) follows from (42). Rearranging the terms completes the proof. ∎
Finally, we note that Theorem 18 makes the claim that $\frac{\alpha_{\theta_{t},\overline{\theta}_{t}}}{\beta_{\theta_{t}}}\in(0,1]$ (using the notation from the same theorem). We now provide the proof for this result.
Proposition C.1 (The RSC to smoothness ratio).
Consider the setting of Theorem 18. Then, $\frac{\alpha_{\theta_{t},\overline{\theta}_{t}}}{\beta_{\theta_{t}}}\in(0,1]$.
Proof.
From Lemma 41 we have
${\cal L}(\theta_{t})-{\cal L}(\bar{\theta}_{t})\leq\frac{1}{2\alpha_{\theta_{t},\overline{\theta}_{t}}}\left\lVert\nabla_{\theta}{\cal L}(\theta_{t})\right\rVert_{2}^{2}.$ (43)
Now, using Lemma 5.2, we obtain
${\cal L}(\theta_{t+1})$ $\leq{\cal L}(\theta_{t})+\langle\theta_{t+1}-\theta_{t},\nabla_{\theta}{\cal L}(\theta_{t})\rangle+\frac{\beta_{\theta_{t}}}{2}\|\theta_{t+1}-\theta_{t}\|_{2}^{2}$ (44)
$={\cal L}(\theta_{t})-\eta_{t}\left\lVert\nabla_{\theta}{\cal L}(\theta_{t})\right\rVert_{2}^{2}+\frac{\beta_{\theta_{t}}}{2}\eta_{t}^{2}\|\nabla_{\theta}{\cal L}(\theta_{t})\|_{2}^{2}$
$={\cal L}(\theta_{t})-\frac{1}{2\beta_{\theta_{t}}}\left\lVert\nabla_{\theta}{\cal L}(\theta_{t})\right\rVert_{2}^{2}$
where the last equality follows from $\eta_{t}=\frac{1}{\beta_{\theta_{t}}}$.
Now, we notice that $\theta_{t+1}\in Q^{\theta_{t}}_{\kappa}$ by Definition 5.1, $\theta_{t+1}\in B_{\rho_{2}}^{\operatorname{Euc}}(\theta_{t})\cap\{\theta\in\mathbb{R}^{p}\,|\,\mathbf{v}\in B_{\rho_{1}}^{\operatorname{Euc}}(\mathbf{v}_{0})\}$ by Assumption 3, and so $\theta_{t+1}\in B_{t}$. This implies that ${\cal L}(\bar{\theta}_{t})\leq{\cal L}(\theta_{t+1})$. Thus, using (44), we obtain
$\frac{1}{2\beta_{\theta_{t}}}\left\lVert\nabla_{\theta}{\cal L}(\theta_{t})\right\rVert_{2}^{2}\leq{\cal L}(\theta_{t})-{\cal L}(\bar{\theta}_{t})~{}.$
Using this inequality along with (43) leads to $\frac{\alpha_{\theta_{t},\overline{\theta}_{t}}}{\beta_{\theta_{t}}}\leq 1$. Finally, by assumption we have $\alpha_{\theta_{t},\overline{\theta}_{t}}>0$ and by (17) we have $\beta_{\theta_{t}}>0$, and so $\frac{\alpha_{\theta_{t},\overline{\theta}_{t}}}{\beta_{\theta_{t}}}>0$. This completes the proof. ∎
C.3 Distance between consecutive iterates
Proposition C.2 (Bound on the distance between consecutive iterates for WeightNorm under $\phi(0)=0$).
Under Assumptions 1 and 2, and assuming (A3.1) from Assumption 3 and that the activation function satisfies $\phi(0)=0$, we have
$\left\lVert\theta_{t+1}-\theta_{t}\right\rVert_{2}\leq\eta_{t}\cdot{\cal O}\left((1+\rho_{1})^{2}\left(1+\frac{\sqrt{L}}{\sqrt{m}\left\lVert\bar{W}_{t}\right\rVert_{2}}\right)\right).$
Proof.
We have that $\left\lVert\theta_{t+1}-\theta_{t}\right\rVert_{2}=\left\lVert\theta_{t}-\eta_{t}\nabla_{\theta}{\cal L}(\theta_{t})-\theta_{t}\right\rVert_{2}=\eta_{t}\left\lVert\nabla_{\theta}{\cal L}(\theta_{t})\right\rVert_{2}\overset{(a)}{\leq}\eta_{t}\cdot 2\varrho_{\theta_{t}}\sqrt{\varphi}\overset{(b)}{\leq}\eta_{t}\cdot{\cal O}\left((1+\rho_{1})\left(1+\frac{\sqrt{L}(1+\rho_{1})}{\sqrt{m}\left\lVert\bar{W}_{t}\right\rVert_{2}}\right)\right)\leq\eta_{t}\cdot{\cal O}\left((1+\rho_{1})^{2}\left(1+\frac{\sqrt{L}}{\sqrt{m}\left\lVert\bar{W}_{t}\right\rVert_{2}}\right)\right)$, where (a) follows from using (10) from Proposition 4.1 and the definitions of $\varrho_{\theta_{t}}$ and $\varphi$ presented therein; and (b) follows from Corollary 4.2 and Corollary 4.3. ∎
Appendix D Comparison between networks with and without WeightNorm
We summarize in Table 1 the comparison between our bounds and optimization results and those for networks without WeightNorm by Banerjee et al. [2023].
with WeightNorm without WeightNorm
Deterministic results High probability results
Hidden layer weights can take any value Hidden layer weights restricted to a neighborhood around their initialization
Appearance of normalization terms, i.e., the minimum weight vector norm, in the Hessian bounds and the RSC and smoothness parameters No appearance of such normalization terms
The polynomial dependence on the depth in our bounds and parameters is independent from the initialization variance of the weights Polynomial dependence on the depth requires careful choice of initialization variance
Lipschitz constant bound is independent from the depth Lipschitz constant bound exponentially grows with the depth unless a careful choice of initialization variance is done
Table 1: Comparison of results for neural networks with and without WeightNorm.
Appendix E Generalization bound for WeightNorm